### v1

# Integrated Sensing and Communications for Pinching-Antenna Systems (PASS)

Abstract
An integrated sensing and communication (ISAC) design for pinching antenna systems (PASS) is proposed, where the pinching antennas are deployed for establishing reliable line-of-sight communication and sensing links. More particularly, a separated ISAC design is proposed for the two-waveguide PASS, where one waveguide is used to emit the joint communication and sensing signals while the other waveguide is used to receive the reflected echo signals. Based on this framework, a penalty-based alternating optimization algorithm is proposed to maximize the illumination power as well as ensure the communication quality-of-service requirement. Numerical results demonstrate that 1) the proposed PASS-ISAC scheme outperforms the other baseline schemes, and 2) the considered equal power allocation model achieves a performance comparable to the optimal power allocation.
Index Terms:
Beamforming design, integrated sensing and communication, pinching antenna systems.
I Introduction
Fuelled by the burgeoning demands for massive data transmission and pervasive network coverage, flexible antennas have emerged as a promising technique for sixth-generation (6G) cellular systems. Benefiting from their ability to reconfigure the wireless channel, flexible antennas can significantly enhance the throughput of wireless networks. However, traditional flexible antennas (e.g., moveable antennas [1] and fluid antennas [2]) merely permit the adjustment of the antenna position within a range of orders of magnitude comparable to the carrier wavelength. Against this backdrop, the pinching antenna has emerged [3], which is a type of dielectric waveguide-based leaky wave antenna. By applying dielectric particles to a particular point on the dielectric waveguide, a pinching antenna can be activated to establish EM radiation fields and form a communication area [4]. Then, the EM signal inside the dielectric waveguide will be radiated from the pinching antenna to free space with a defined phase shift adjustment (referred to as the pinching beamformer). Notably, as the dielectric waveguide can be pinched at any position to radiate radio waves, the pinching antenna can flexibly move along the dielectric waveguide over a length of dozens of meters, thereby relocating to the closest position to the receiver and creating reliable LoS links.
To enable emerging applications, such as autonomous driving, extended reality, and Metaverse, sensing functionality is recognized as an important indicator of future networks. In pursuit of this vision, the integrated sensing and communication (ISAC) technology has drawn significant attention recently [5], which aims to leverage the cellular network hardware platforms and dedicated signal processing algorithms to achieve the incorporation of communication and sensing functionalities. Recently, it has been claimed that conducting ISAC transmission in the pinching antenna systems (PASS) can further upgrade the communication and sensing (C&S) performance of the network [6]. On the one hand, the pinching antenna can be flexibly repositioned to augment the echo signal energy. On the other hand, the wide-range mobility characteristic of pinching antennas results in an antenna aperture spanning dozens of meters. It inherently enables near-field sensing, e.g., the possibility of simultaneous angular and distance information estimation and even target velocity sensing, thereby offering a more comprehensive and accurate sensing of the surrounding environment. Nevertheless, as of the present moment, research in the PASS-ISAC remains conspicuously absent.
Motivated by the above, this paper proposes a separated ISAC design for PASS. To elaborate, the BS is connected with two dielectric waveguides, where one waveguide is used to transmit the communication and sensing signals, while the other is employed to collect the reflected echo signals from the target. We aim to maximize the illumination power at the target while satisfying the quality-of-service (QoS) requirement of the communication user by optimizing the pinching beamforming offered by the mobility of pinching antennas. A penalty-based alternating optimization (AO) algorithm is proposed to handle the non-convex optimization problem, where the optimal radiation pattern is obtained under an arbitrary antenna activation configuration. Numerical results evaluate the superiority of the proposed scheme over the baseline schemes. It also reveals that the equal power model achieves a performance comparable to the optimal power allocation.
II System Model and Problem Formulation
Figure 1: The separated ISAC design for PASS.
As shown in Fig. 1, we consider a PASS-ISAC system, where a dual-function base station (BS) conveys with a single-antenna communicating user while sensing a point-like target. The BS is connected with two dielectric waveguides, where one waveguide is used to emit the downlink signals, and the other is employed to receive the reflected echo signals from the target. We assume each dielectric waveguide consists of $N$ pinching antennas. To support simultaneous C&S transmission with the independent pinching beamformer design, we propose a functionality-split structure for PASS, where the transmitting antenna array is partitioned into two portions. To elaborate, the former $N_{1}$ pinching antennas serve for emitting the information-bearing signals to the communication user (referred to as the communication array), and the latter $N_{2}=N-N_{1}$ antennas (referred to as the sensing array) used to sense the target. Then, the reflected echoes from the target would be collected at the receiving antennas, which are transmitted to the BS for parameter estimation.
A three-dimensional (3D) coordination system is considered, where two dielectric waveguides extended from the BS are assumed to be parallel to the x-axis with respect to the x-o-y plane at a distance $d$. The position of the $n$th pinching antenna distributed along the transmitting and receiving dielectric waveguides can be denoted as $\psi_{n}^{\text{p}}=(x_{n}^{\text{p}},0,d)$ and $\psi_{n}^{\text{q}}=(x_{n}^{\text{q}},\bar{y},d)$. The communication user and sensing target are located in the x-o-y plane, with the coordinates of $\psi^{\text{c}}=(x_{1},y_{1},0)$ and $\psi^{\text{s}}=(x_{2},y_{2},0)$, respectively. Furthermore, we assume the target is a static node or moves at a low speed. Thus, the Doppler effect is neglected in this work.
In the considered network, the pinching antennas are non-uniformly disposed on the dielectric waveguide covering the entire range of the user’s activity, which implies that the aperture of pinching antennas may have the same order of magnitude as the signal transmission distance. Without loss of accuracy, we adopt the spherical-wave-based near-field channel model, where only the LoS path is considered. Let $r_{\text{s}}$ and $\phi_{\text{s}}$ denote the distance of the target relative to the origin of the coordinate system and the azimuth, respectively. The coordinate of the target can be rewritten as $(r_{\text{s}}\cos\phi_{\text{s}},r_{\text{s}}\sin\phi_{\text{s}},0)$. Consequently, the distance from the $n$-th pinching antenna to the target is given by
$r_{n}(r_{\text{s}},\phi_{\text{s}})$ $=\|\psi^{\text{s}}-\psi_{n}^{\text{p}}\|$
$=\sqrt{r_{\text{s}}^{2}-2r_{\text{s}}\cos\phi_{\text{s}}x_{n}^{\text{p}}+(x_{n}^{\text{p}})^{2}+d^{2}}.$ (1)
Thus, the channel vector from the transmitting antenna arrays to the target and the communication user can be expressed as
$\mathbf{h}_{\text{s}}(r_{\text{s}},\phi_{\text{s}})=\left[\frac{\eta^{\frac{1}{2}}e^{\jmath\frac{2\pi}{\lambda}r_{1}(r_{\text{s}},\phi_{\text{s}})}}{r_{1}(r_{\text{s}},\phi_{\text{s}})},\cdots,\frac{\eta^{\frac{1}{2}}e^{\jmath\frac{2\pi}{\lambda}r_{N}(r_{\text{s}},\phi_{\text{s}})}}{r_{N}(r_{\text{s}},\phi_{\text{s}})}\right]^{T},$ (2)
$\mathbf{h}_{\text{c}}=\left[\frac{\eta^{\frac{1}{2}}e^{\jmath\frac{2\pi}{\lambda}\|\psi^{\text{c}}-\psi_{1}^{\text{p}}\|}}{\|\psi^{\text{c}}-\psi_{1}^{\text{p}}\|},\cdots,\frac{\eta^{\frac{1}{2}}e^{\jmath\frac{2\pi}{\lambda}\|\psi^{\text{c}}-\psi_{N}^{\text{p}}\|}}{\|\psi^{\text{c}}-\psi_{N}^{\text{p}}\|}\right]^{T},$ (3)
where $\lambda=\frac{2\pi}{f_{\text{c}}}$ dentes the wavelength, $f_{\text{c}}$ is the frequency of the carrier wave, $\eta=\frac{c^{2}}{16\pi^{2}f_{\text{c}}^{2}}$, and $c$ denotes the speed of light.
II-A ISAC Model for Pinching Antenna Networks
Consider a coherent time block of length $T$, the BS transmits the superimposed communication and sensing waveforms to the dielectric waveguide. We assume the communication channel condition and the sensing parameters remain unchanged during one coherent time block. Thus, the emitted signal at the $t$-th time slot is given by $s(t)\in\mathbb{C}$, which is assumed to be normalized and independently distributed, i.e., $\mathbb{E}\{|s(t)|^{2}\}=1$ and $\mathbb{E}\{s(t)s^{*}(\bar{t})\}=0$. On receiving $s(t)$, the dielectric waveguide radiates the signal using the pinching beamformer. The radiated signal from the pinching antenna is given by
$\mathbf{x}(t)=\left[\sqrt{P_{1}}e^{\jmath\theta_{1}},\cdots,\sqrt{P_{N}}e^{\jmath\theta_{N}}\right]s(t),$ (4)
where $P_{n}$ and $\theta_{n}$ denote the power allocation and the radiation phase shift at the $n$-th pinching antenna. For ease of implementation, the equal power allocation model is assumed, i.e., $P_{1}=\cdots=P_{N}=\frac{P_{\text{max}}}{N}$ [4]. $\theta_{n}$ is defined by $2\pi\eta_{\text{eff}}\frac{\|\psi_{0}^{\text{p}}-\psi_{n}^{\text{p}}\|}{\lambda}$, where $\psi_{0}^{\text{p}}$ denotes the location of the feed point, and $\eta_{\text{eff}}$ denotes the effective refractive index of the dielectric waveguide.
To facilitate the beamforming design, we consider a semi-continuous activation structure for pinching antennas. As shown in Fig. 2, the pinching antennas can be activated in a large-scale range over the dielectric waveguide to change the free-space path loss and the spherical-wave array response. However, when the large-scale activation position is determined, the pinching antennas can also be adjusted in a small-scale range to change the phase shift of the radiation signal, but without significantly affecting the wireless channel characteristics. For example, consider a millimeter wave ISAC scenario with a carrier frequency of $28$ GHz, a dielectric waveguide length of 50 meters (m), and $\eta_{\text{eff}}=1.4$, the small-scale phase-shift adjustment range will not exceed a maximum of $7.7$ millimeter, which has a negligible impact on the large-scale path loss as well as the spherical-wave array response. Consequently, it can be reasonably assumed that the radiation phase shift design of pinching antennas is independent of the large-scale activation position optimization.
Figure 2: Semi-continuous activation structure.
In this paper, we aim to utilize the communication signal to achieve simultaneous communication and target sensing. Thus, the transmitted signal can be rewritten as $\mathbf{x}(t)=[\mathbf{w}s(t),\mathbf{v}s(t)]=[\mathbf{w}s(t),\mathbf{s}(t)]$, where $\mathbf{w}s(t)$ is the communication signal and $\mathbf{v}s(t)=\mathbf{s}(t)$ represents the equivalent sensing signal. Here, we have $\mathbf{w}=\left[\sqrt{P_{1}\alpha_{\text{c}}}e^{\jmath\theta_{1}},\cdots,\sqrt{P_{N_{1}}\alpha_{\text{c}}}e^{\jmath\theta_{N_{1}}}\right]$, and $\mathbf{v}=\left[\sqrt{P_{N_{1}+1}\alpha_{\text{c}}}e^{\jmath\theta_{N_{1}+1}},\cdots,\sqrt{P_{N}\alpha_{\text{c}}}e^{\jmath\theta_{N}}\right]$.
II-A1 Communication Performance Metric
With the aforementioned signal model, the received signals at the communication user are given by
$y(t)=\mathbf{h}_{[1:N_{1}],\text{c}}^{H}\mathbf{w}s(t)+\mathbf{h}_{[N_{1}+1:N],\text{c}}^{H}\mathbf{s}(t)+n(t),$ (5)
where $n(t)\sim\mathcal{CN}(0,\sigma^{2})$ denotes the additive white Gaussian noise (AWGN) at the communication user, and $\mathbf{h}_{[m:n],\xi}$ represents a subvector of $\mathbf{h}_{\xi}$ consisting of the elements with indexes from $m$ to $n$, $(m<n,\xi\in\{\text{c},\text{s}\})$. Hence, the achievable rate of the communication user is given by
$R=\log_{2}\left(1+\frac{|\mathbf{h}_{1,\text{c}}^{H}\mathbf{w}|^{2}}{\mathbf{h}_{2,\text{c}}^{H}\mathbf{R}_{\text{s}}\mathbf{h}_{2,\text{c}}^{\text{c}}+\sigma^{2}}\right),$ (6)
where $\mathbf{R}_{\text{s}}=\mathbf{v}\mathbf{v}^{H}$. For simplicity of expression, we abbreviate $\mathbf{h}_{[1:N_{1}],\text{c}}$ and $\mathbf{h}_{[N_{1}+1:N],\text{c}}$ to $\mathbf{h}_{1,\text{c}}$ and $\mathbf{h}_{2,\text{c}}$ in the following.
II-A2 Sensing Performance Metric
For target sensing, we adopt the illumination power as the performance metric, which characterizes the received sensing signal power at the target [7]. Thus, the illumination power with respect to azimuth angle $\phi_{\text{s}}$ and distance $r_{\text{s}}$ is given by
$P(\theta_{\text{s}},r_{\text{s}})$ $=\mathbb{E}\left\{\left|\mathbf{h}_{[1:N_{1}],\text{s}}^{H}(r_{\text{s}},\phi_{\text{s}})\mathbf{w}s(t)\right|^{2}\right\}+$
$\mathbb{E}\left\{\left|\mathbf{h}_{[N_{1}+1:N],\text{s}}^{H}(r_{\text{s}},\phi_{\text{s}})\mathbf{s}(t)\right|^{2}\right\}$
$=\mathbf{h}_{1,\text{s}}^{H}\mathbf{w}\mathbf{w}^{H}\mathbf{h}_{1,\text{s}}+\mathbf{h}_{2,\text{s}}^{H}\mathbf{R}_{\text{s}}\mathbf{h}_{2,\text{s}},$ (7)
where $\mathbf{h}_{[1:N_{1}],\text{s}}(r_{\text{s}},\phi_{\text{s}})=\mathbf{h}_{1,\text{s}}$ and $\mathbf{h}_{[N_{1}+1:N],\text{s}}(r_{\text{s}},\phi_{\text{s}})=\mathbf{h}_{2,\text{s}}$.
II-B Problem Formulation
In this paper, we aim to maximize the illumination power $P(\theta_{\text{s}},r_{\text{s}})$ by jointly designing the pinching beamformer and antenna-activation positions, under the transmit power budget and communication QoS requirement, which is given by
$(\text{P}1)\quad$ $\max\limits_{\mathbf{h}_{1,\text{s}},\mathbf{h}_{2,\text{s}},\mathbf{h}_{1,\text{c}},\atop\mathbf{h}_{2,\text{c}},\mathbf{w},\mathbf{R}_{\text{s}}}P(\theta_{\text{s}},r_{\text{s}})$ (8a)
s.t. $|\mathbf{w}_{[n]}|^{2}=\mathbf{R}_{[n,n],\text{s}}=\frac{P_{\text{max}}}{N},$ (8b)
$\mathbf{R}\geq R_{\text{QoS}},$ (8c)
$\mathbf{R}_{\text{s}}\succeq\mathbf{0},\ \text{rank}(\mathbf{R}_{\text{s}})=1,$ (8d)
$\mathbf{h}_{[n],\text{s}}=\mathcal{H}_{\text{s}}(n),$ (8e)
$\mathbf{h}_{[n],\text{c}}=\mathcal{H}_{\text{c}}(n),$ (8f)
$|x_{n}^{\text{p}}-x_{m}^{\text{p}}|\geq\Delta x,\ n\neq m,$ (8g)
$x_{n}^{\text{p}}\in[x_{n}^{\text{p,min}},x_{n}^{\text{p,max}}],\ n\in[1,N],$ (8h)
where $\mathcal{H}_{\text{s}}(n)=\frac{\eta^{\frac{1}{2}}e^{\jmath\frac{2\pi}{\lambda}\sqrt{r_{\text{s}}^{2}-2r_{\text{s}}\cos\phi_{\text{s}}x_{n}^{\text{p}}+(x_{n}^{\text{p}})^{2}+d^{2}}}}{\sqrt{r_{\text{s}}^{2}-2r_{\text{s}}\cos\phi_{\text{s}}x_{n}^{\text{p}}+(x_{n}^{\text{p}})^{2}+d^{2}}}$, $\mathcal{H}_{\text{c}}(n)=\frac{\eta^{\frac{1}{2}}e^{\jmath\frac{2\pi}{\lambda}\sqrt{x_{1}^{2}-2x_{1}x_{n}^{\text{p}}+(x_{n}^{\text{p}})^{2}+y_{1}^{2}+d^{2}}}}{\sqrt{x_{1}^{2}-2x_{1}x_{n}^{\text{p}}+(x_{n}^{\text{p}})^{2}+y_{1}^{2}+d^{2}}}$, $P_{\text{max}}$ denotes the maximal transmit power at the BS, and $R_{\text{QoS}}$ denotes the minimum tolerated achievable rate of the communication user. (8b) denotes the maximal transmit power constraint; (8c) represents the QoS requirement; (8g) and (8h) limits the activation range of each pinching antenna. Notably, the problem (P1) is challenging to solve due to the quadratic objective function and the coupling between the $\{\mathbf{h}_{1,\text{s}},\mathbf{h}_{2,\text{s}},\mathbf{h}_{1,\text{c}},\mathbf{h}_{2,\text{c}}\}$ and $\{\mathbf{w},\mathbf{R}_{\text{s}}\}$.
III Pinching Beamforming Optimization
In this section, we focus on the pinching beamforming design by jointly optimizing the radiation pattern and activation position of pinching antennas. Since all the channels $\{\mathbf{h}_{1,\text{s}},\mathbf{h}_{2,\text{s}},\mathbf{h}_{1,\text{c}},\mathbf{h}_{2,\text{c}}\}$ are relative to $\mathbf{x}^{\text{p}}$, which denotes the x-axis coordinate of the transmitting pinching antennas. Thus, the problem (P1) can be rewritten as
$(\text{P}2)\ \max\limits_{\mathbf{x}^{\text{p}},\mathbf{w},\mathbf{R}_{\text{s}}}\quad$ $\mathbf{h}_{1,\text{s}}^{H}\mathbf{w}\mathbf{w}^{H}\mathbf{h}_{1,\text{s}}+\mathbf{h}_{2,\text{s}}^{H}\mathbf{R}_{\text{s}}\mathbf{h}_{2,\text{s}}$ (9a)
s.t. $|(\mathbf{h}_{1,\text{c}})^{H}\mathbf{w}|^{2}\geq\gamma_{\text{QoS}}[(\mathbf{h}_{2,\text{c}})^{H}\mathbf{R}_{\text{s}}\mathbf{h}_{2,\text{c}}+\sigma^{2}],$ (9b)
[…]
where $\gamma_{\text{QoS}}=2^{R_{\text{QoS}}}-1$. To address the quadratic objective and constraints, we apply the SDR technique to rewrite the problem (P2) as follows.
$(\text{P}3)\ \max\limits_{\mathbf{x}^{\text{p}},\mathbf{W},\mathbf{R}_{\text{s}}}$ $\text{Tr}(\mathbf{H}_{1,\text{s}}\mathbf{W})+\text{Tr}(\mathbf{H}_{2,\text{s}}\mathbf{R}_{\text{s}})$ (10a)
s.t. $\mathbf{W}_{[n,n]}=\mathbf{R}_{[n,n],\text{s}}=\frac{P_{\text{max}}}{N},$ (10b)
$\text{Tr}(\mathbf{H}_{1,\text{c}}\mathbf{W})\geq\gamma_{\text{QoS}}[\text{Tr}(\mathbf{H}_{2,\text{c}}\mathbf{R}_{\text{s}})+\sigma^{2}],$ (10c)
$\mathbf{W}\succeq\mathbf{0},\ \mathbf{R}_{\text{s}}\succeq\mathbf{0},$ (10d)
$\text{rank}(\mathbf{W})=1,\ \text{rank}(\mathbf{R}_{\text{s}})=1,$ (10e)
[…]
where $\mathbf{W}=\mathbf{w}\mathbf{w}^{H}$, $\mathbf{H}_{1,\text{s}}=\mathbf{h}_{1,\text{s}}\mathbf{h}_{1,\text{s}}^{H}$, $\mathbf{H}_{2,\text{s}}=\mathbf{h}_{2,\text{s}}\mathbf{h}_{2,\text{s}}^{H}$, $\mathbf{H}_{1,\text{c}}=\mathbf{h}_{1,\text{c}}\mathbf{h}_{1,\text{c}}^{H}$, and $\mathbf{H}_{2,\text{c}}=\mathbf{h}_{2,\text{c}}\mathbf{h}_{2,\text{c}}^{H}$. In order to deal with the coupled optimization variables $\{\mathbf{x}^{\text{p}}\}$ and $\{\mathbf{W},\mathbf{R}_{\text{s}}\}$, we consider the penalty-based AO framework below.
III-A Radiation Pattern Design
With the given $\{\mathbf{x}^{\text{p}}\}$, the problem (P3) is reformulated as
$(\text{P}4)\ \max\limits_{\mathbf{W},\mathbf{R}_{\text{s}}}\quad$ $\text{Tr}(\mathbf{H}_{1,\text{s}}\mathbf{W})+\text{Tr}(\mathbf{H}_{2,\text{s}}\mathbf{R}_{\text{s}})$ (11a)
[…]
The problem (P4) is a SDP problem with rank-one constraints, which contain $3$ feasible linear constraints and $2$ matrix variables. According to [8, Eq. (32)], it readily knows the optimal general-rank solutions of (P4) satisfy $\text{rank}(\mathbf{W})^{2}+\text{rank}(\mathbf{R}_{\text{s}})^{2}\leq 3$, which indicates that $\text{rank}(\mathbf{W})=\text{rank}(\mathbf{R}_{\text{s}})=1$. By applying the conclusion, the optimal pinching beamformer can be obtained.
III-B Antenna Activation Position Optimization
To deal with the intractable expression of variable channels, we introduce auxiliary variables $\mathbf{U}_{1}$, $\mathbf{V}_{1}$, $\mathbf{U}_{2}$, and $\mathbf{V}_{2}$ to replace $\mathbf{H}_{1,\text{s}}$, $\mathbf{H}_{1,\text{c}}$, $\mathbf{H}_{2,\text{s}}$, and $\mathbf{H}_{2,\text{c}}$, respectively. Thus, we have the equality constraints $\mathbf{U}_{1}=\mathbf{H}_{1,\text{s}}$, $\mathbf{V}_{1}=\mathbf{H}_{2,\text{s}}$, $\mathbf{U}_{2}=\mathbf{H}_{1,\text{c}}$, and $\mathbf{V}_{2}=\mathbf{H}_{2,\text{c}}$. By relocating the equality constraint to the objective function and serving as a penalty term, the problem (P3) can be equivalently rewritten as
$(\text{P}5)\quad$ $\max\limits_{\mathbf{x}^{\text{p}},\mathbf{U}_{i},\mathbf{V}_{i}}\text{Tr}(\mathbf{U}_{1}\mathbf{W})+\text{Tr}(\mathbf{V}_{1}\mathbf{R}_{\text{s}})-\frac{1}{2\varrho}\chi$ (12a)
s.t. $\text{Tr}(\mathbf{U}_{2}\mathbf{W})\geq\gamma_{\text{QoS}}[\text{Tr}(\mathbf{V}_{2}\mathbf{R}_{\text{s}})+\sigma^{2}],$ (12b)
$\text{rank}(\mathbf{U})=1,\ \text{rank}(\mathbf{V})=1,$ (12c)
$\mathbf{U}\succeq\mathbf{0},\ \mathbf{V}\succeq\mathbf{0},$ (12d)
$\frac{1}{r_{\text{min,s}}^{2}}\leq\frac{\mathbf{X}_{[n,n]}}{\eta}\leq\frac{1}{r_{\text{max,s}}^{2}},\ \mathbf{X}\in\{\mathbf{U}_{1},\mathbf{V}_{1}\},$ (12e)
$\frac{1}{r_{\text{min,c}}^{2}}\leq\frac{\mathbf{Y}_{[n,n]}}{\eta}\leq\frac{1}{r_{\text{max,c}}^{2}},\ \mathbf{Y}\in\{\mathbf{U}_{2},\mathbf{V}_{2}\},$ (12f)
[…]
where $\chi=\|\mathbf{U}_{1}-\mathbf{H}_{1,\text{s}}\|_{F}+\|\mathbf{V}_{1}-\mathbf{H}_{2,\text{s}}\|_{F}+\|\mathbf{U}_{2}-\mathbf{H}_{1,\text{c}}\|_{F}+\|\mathbf{V}_{2}-\mathbf{H}_{2,\text{c}}\|_{F}$, and $\varrho$ denotes the scaling factor of the penalty terms. Here, $r_{\text{min,s}}$ and $r_{\text{max,s}}$ denote the minimum and maximum distances from an arbitrary pinching antenna to the sensing target. Also, $r_{\text{min,c}}$ and $r_{\text{max,c}}$ denote the minimum and maximum distances from an arbitrary pinching antenna to the communication user. To solve the problem (P5), we propose a penalty-based two-layer algorithm, which alternately optimizes $\{\mathbf{U},\mathbf{V}\}$ and $\{\mathbf{x}^{\text{p}}\}$ in the inner layer and update $\varrho$ in the outer layer.
III-B1 Inner layer iteration—subproblem with respect to $\{\mathbf{U},\mathbf{V}\}$
With the fixed $\{\mathbf{x}^{\text{p}}\}$, the problem (P5) is reduced to
$(\text{P}6)\ \max\limits_{\mathbf{x}^{\text{p}},\mathbf{U}_{i},\mathbf{V}_{i}\atop\mathbf{W},\mathbf{R}_{\text{s}}}$ $\text{Tr}(\mathbf{U}_{1}\mathbf{W})+\text{Tr}(\mathbf{V}_{1}\mathbf{R}_{\text{s}})-\frac{1}{2\varrho}\chi$ (13a)
[…]
To handle the rank-one constraint, we employ the difference-of-convex (DC) relaxation method [9] to rewrite the (12c) as
$\begin{cases}\Re(\text{Tr}(\mathbf{U}_{i}^{H}(\mathbf{I}-\mathbf{u}_{i}\mathbf{u}_{i}^{H})))\leq\varrho_{1},\\ \Re(\text{Tr}(\mathbf{V}_{i}^{H}(\mathbf{I}-\mathbf{v}_{i}\mathbf{v}_{i}^{H})))\leq\varrho_{2},\end{cases}\ i\in\{1,2\},$ (14)
where $\mathbf{u}_{i}$ and $\mathbf{v}_{i}$ represent the leading eigenvectors of $\mathbf{U}_{i}$ and $\mathbf{V}_{i}$. As a result, the problem (P6) can be transformed into
$(\text{P}7)\ \max\limits_{\mathbf{U}_{i},\mathbf{V}_{i},\varrho_{i}}\quad$ $\text{Tr}(\mathbf{U}_{1}\mathbf{W})+\text{Tr}(\mathbf{V}_{1}\mathbf{R}_{\text{s}})-\frac{1}{2\varrho}\chi$ (15a)
s.t. $\varrho_{i}>0,\ i\in\{1,2\},$ (15b)
[…]
which is a convex problem and can be directly solved. Thus, the rank-one solution $\{\mathbf{U}_{i},\mathbf{V}_{i}\}$ can be obtained by iteratively solving the problem (P7).
III-B2 Inner layer iteration—subproblem with respect to $\{\mathbf{x}^{\text{p}}\}$
Note that the equality constraint $\mathbf{U}_{1}=\mathbf{H}_{1,\text{s}}$, $\mathbf{V}_{1}=\mathbf{H}_{2,\text{s}}$, $\mathbf{U}_{2}=\mathbf{H}_{1,\text{c}}$ are equivalent to $\mathbf{u}_{1}=\mathbf{h}_{1,\text{s}}$, $\mathbf{v}_{1}=\mathbf{h}_{2,\text{s}}$, $\mathbf{u}_{2}=\mathbf{h}_{1,\text{c}}$ and $\mathbf{v}_{2}=\mathbf{h}_{2,\text{c}}$, where $\{\mathbf{u}_{1},\mathbf{v}_{1},\mathbf{u}_{2},\mathbf{v}_{2}\}$ denote the leading eigenvectors of $\{\mathbf{U}_{1},\mathbf{V}_{1},\mathbf{U}_{2},\mathbf{V}_{2}\}$. As a result, the problem (P6) can be transformed into
$(\text{P}8)\ \min\limits_{\mathbf{x}^{\text{p}}}\quad$ $\|\mathbf{u}_{1}-\mathbf{h}_{1,\text{s}}\|^{2}+\|\mathbf{v}_{1}-\mathbf{h}_{2,\text{s}}\|^{2}+\|\mathbf{u}_{2}-\mathbf{h}_{1,\text{c}}\|^{2}$
$+\|\mathbf{v}_{2}-\mathbf{h}_{2,\text{c}}\|^{2}$ (16a)
[…]
It is easy to notice that $x_{n}^{\text{p}}$ and $x_{m}^{\text{p}}$ ($n\neq m$) are separated in the objective function but coupled in the constraint (8h), which motivates us to adopt the element-wise optimization framework. Therefore, with the fixed $\{x_{1}^{\text{p}},\cdots,x_{n-1}^{\text{p}},x_{n+1}^{\text{p}},\cdots,x_{N}^{\text{p}}\}$, the subproblem with respect to $x_{n}^{\text{p}}$ is given by
$(\text{P}9)\ \min\limits_{x_{n}^{\text{p}}}\quad$ $\varpi(x_{n}^{\text{p}})$ (17a)
s.t. $x_{n}^{\text{p}}\in[x_{n-1}^{\text{p}}+\Delta x,x_{n+1}^{\text{p}}-\Delta x],$ (17b)
where $\varpi(x_{n}^{\text{p}})$ is given by
$\varpi(x_{n}^{\text{p}})=\begin{cases}\|\mathbf{u}_{[n],1}-\mathbf{h}_{[n],\text{s}}\|^{2}+\|\mathbf{u}_{[n],2}-\mathbf{h}_{[n],\text{c}}\|^{2},\ n\in\mathcal{N}_{1},\\ \|\mathbf{v}_{[n],1}-\mathbf{h}_{[n],\text{s}}\|^{2}+\|\mathbf{v}_{[n],2}-\mathbf{h}_{[n],\text{c}}\|^{2},\ n\in\mathcal{N}_{2}.\end{cases}$ (18)
where $\mathcal{N}_{1}=\{1,\cdots,N_{1}\}$ and $\mathcal{N}_{2}=\{N_{1}+1,\cdots,N\}$. Then, the optimal $x_{n}^{\text{p}}$ can be obtained by the low-complexity one-dimensional search..
III-B3 Outer layer iteration
In the outer layer, we initialise a large $\varrho$ and update $\varrho$ at each outer iteration by $\varrho=\varrho\bar{c}$, where $10<\bar{c}<$ 1 is the iteration coefficient of the penalty terms. The penalty-based two-layer algorithm is summarized in Algorithm 1.
III-C Overall Algorithm
Algorithm 1 Penalty-based two-layer algorithm.
1:  Initialize $\mathbf{x}^{\text{p}}$, $\mathbf{u}_{i}$, and $\mathbf{v}_{i}$. Set the convergence accuracy $\epsilon_{1}$, $\epsilon_{2}$, and $\epsilon_{3}$.
2:  repeat
3:     repeat
4:         update $\{\mathbf{U},\mathbf{V}\}$ by iteratively solving the subproblem (P7) with an accuracy of $\epsilon_{1}$.
5:         update $\mathbf{x}^{\text{p}}$ by adopting element-wise exhaustive search.
6:     until  the objective value converges with an accuracy of $\epsilon_{2}$.
7:      $\varrho=\varrho\bar{c}$.
8:  until  $\varrho\leq\epsilon_{3}$.
Algorithm 2 Penalty-based AO algorithm.
1:  Parameter Initialization. Set a convergence accuracy $\epsilon_{4}$.
2:  repeat
3:      update $\{\mathbf{W},\mathbf{R_{\text{s}}}\}$ by solving the problem (P4).
4:      carry out Algorithm 1.
5:  until the objective value converges with an accuracy of $\epsilon_{4}$.
The proposed penalty-based AO algorithm is summarized in Algorithm 2, which is assured to converge at least to a stationary point solution. The computational complexity of Algorithm 2 mainly depends on solving the SDP problems (P4), (P7), and the one-dimensional exhaustive search. It is given by $\mathcal{O}\Big{(}\log(\frac{1}{\epsilon_{4}})((N_{1})^{3.5}+\log(\frac{1}{\epsilon_{3}})\log(\frac{1}{\epsilon_{2}})\big{[}\log(\frac{1}{\epsilon_{1}})(N_{1})^{3.5}+N\bar{Q}\big{]})\Big{)}$ [8], where $\bar{Q}$ represents the number of the quantization bits during the one-dimensional exhaustive search.
IV Numerical Results
Figure 3: The equivalent illumination power versus the transmit power at the BS.
Figure 4: The equivalent illumination power versus the rotation angle of the dielectric waveguide..
Figure 5: The equivalent illumination power versus the rotation angle of the dielectric waveguide.
This section evaluates the performance of the proposed PASS-ISAC framework. A 3D topological network setup is considered, where the dielectric waveguide is located in the x-o-z plane with a height of $d$ and a length of $50$ m. The communicating user and the sensing target are located in a square region centered at the origin in the x-o-y plane. The default simulation parameters are set as: $\sigma^{2}=-105$ dBm, $f=28$ GHz, $d=10$ m, $\Delta x=\frac{\lambda}{2}$, $r_{\text{s}}=30$ m, $\phi_{\text{s}}=\frac{\pi}{3}$, $[x_{1},y_{1},z_{1}]=[-15,-15,0]$ m, $N=12$, $\eta_{\text{eff}}=1.4$, $R_{\text{QoS}}=2$ bps/Hz, $\epsilon_{1}=\epsilon_{2}=\epsilon_{3}=\epsilon_{4}=10^{-3}$. The other network parameters are shown in the captions of the figures.
To validate the performance of the proposed scheme, the following baseline schemes are considered in this paper:
  • Conventional antenna: In this scheme, we deploy $N$ conventional uniform linear array (ULA) at the BS, where the antenna space is set as $\frac{\lambda}{2}$.
  • Fixed antenna: In this scheme, $N$ pinching antennas are uniformly spread along the dielectric waveguide.
  • MRC scheme: In the maximum-ratio combining (MRC) scheme, the communication and sensing beamformer are designed by $\mathbf{w}=\frac{\sqrt{P_{\text{max}}}}{2}\frac{\mathbf{h}_{1,\text{c}}}{\|\mathbf{h}_{1,\text{c}}\|}$ and $\mathbf{s}=\frac{\sqrt{P_{\text{max}}}}{2}\frac{\mathbf{h}_{2,\text{s}}}{\|\mathbf{h}_{s,\text{s}}\|}$ [4].
  • ZF scheme: In the zero-forcing (ZF) scheme, the communication and sensing beamformer are designed by $[\mathbf{w},\mathbf{s}]=\mathbf{H}(\mathbf{H}^{H}\mathbf{H})^{-1}\mathbf{P}$, where $\mathbf{H}=[\mathbf{h}_{1,\text{c}},\mathbf{h}_{2,\text{s}}]$ and $\mathbf{P}=\text{diag}([\frac{\sqrt{P_{\text{max}}}}{2},\frac{\sqrt{P_{\text{max}}}}{2}])$ [4].
Here, the equivalent illumination power is defined as the $\frac{P(\theta_{\text{s}},r_{\text{s}})}{\eta}$, which neglects the impact of the free-space path-loss coefficient $\eta$ on the illumination power. In Fig. 5, we can observe that the pinching antenna achieves the highest illumination power compared to the other baseline schemes. This result can be expected because compared with the conventional antenna and fixed antenna, pinching antennas can be flexibly repositioned to attenuate the large-scale path loss between the pinching antennas and the receiving ends. It indicates that the pinching antenna is capable of providing more spatial degrees-of-freedom (DoFs) to favor the communication and sensing performance. On the other hand, as the MRC and ZF schemes cannot guarantee the optimal beamforming design under the given antenna activation positions, they exhibit inferior performance to the pinching antenna scheme.
Fig. 5 depicts the relationship between the equivalent illumination power and the number of activated pinching antennas, with a comparison of the optimal power allocation model (i.e., each pinching antenna can radiate an arbitrary proportion of the entire transmit power) and proportional power allocation $(P_{n}=\delta(\sqrt{1-\delta^{2}})^{n-1}P_{\text{max}})$ models [10] is considered. Without loss of generality, $\delta$ is set as $0.5$. As can be observed, the equal power allocation model achieves a comparable performance to that of the considered ideal power allocation model. This is because, in the equal power model, each pinching antenna will be allocated a constant level of power, which ensures the spatial DoF gain of the pinching antennas. However, in the proportional power model, pinching antennas located far away from the BS cannot obtain enough power to serve the nearby users, which limits the number of effective pinching antennas. We also observe that increasing $r_{\text{s}}$ decreases the illumination power at the target due to a significant increase in large-scale path loss.
Fig. 5 investigates the impact of the rotation angle of the dielectric waveguide on illumination power at the target. Here, we assume the dielectric waveguide can be rotated in a clockwise direction parallel to the x-o-y plane, where the rotation angle is defined as the angle entwined by the dielectric waveguide and the x-axis. From Fig. 5, it is shown that the illumination power first increases and then decreases as the rotation angle grows. This is due to the fact that when the rotation angle is $60^{\circ}$, the target is located underneath the dielectric waveguide, and it receives the maximum illumination power. As the rotation angle further rises, the distance between the target and the pinching antenna becomes large, so the illumination power gradually decreases. In addition, raising the height of the dielectric waveguide increases the average distance from the pinching antennas to the user and target, thus, the illumination power decreases as $d$ increases.
V Conclusion
A novel PASS-ISAC framework has been proposed, where the transmitting array was evenly split into two parts for supporting independent communication and sensing beamforming designs. A penalty-based AO algorithm was proposed to maximize the illumination power at the target while guaranteeing the QoS requirement of the communication user. Simulation results were provided to verify the superiority of the proposed PASS-ISAC framework over the other benchmarks.